OpenStudy (anonymous):

PLZ HELP PLZZZZZZZZZZZZZZZZZZZZZZZZZZZZZZZ Solve 196⋅x2 - 1,225 = 0 by using square roots

OpenStudy (kirbykirby):

\[ 196x^2-1225=0\\ 196x^2=1225\\ x^2=\frac{1225}{196}\\ x^2=\frac{25}{4}\\ x=\pm \sqrt{\frac{25}{4}}\\ x=\pm\frac{5}{2}\]